OpenStudy (rissyroo118):

What is the slope of a line parallel to the line whose equation is 5x – 3y = 1? help a 1/3 b 5/3 c-3/5 d 3/5

OpenStudy (rissyroo118):

someone please help

OpenStudy (imstuck):

What is the rule about slopes and parallel lines? Do you know it?

OpenStudy (rissyroo118):

no , they didnt give any other info. im sorry

OpenStudy (imstuck):

First to begin, solve this for y. You have to get y alone to get the actual slope of this line first. Ok. For 2 lines to be parallel, they have to have the same slope. That is, they go up or down at the same degree. they never intersect, ever. That's what parallel lines don't ever do, right? So solving the first equation, we get that y = 5/3x-1/3. 5/3 is the slope of the line. It's b in your answer choices there. Always do this: solve for y. Get it all by itself and you will automatically have put the equation into slope-intercept form. "m" is slope and "b" is y-intercept.

OpenStudy (rissyroo118):

okay that kinda makes sense. do you mind helping with a few more problems ?

OpenStudy (imstuck):

I do not mind at all!

OpenStudy (rissyroo118):

thank you! youre awesome

OpenStudy (rissyroo118):

Which is the equation of a line perpendicular to the line with the equation 3x + 4y = 7? a 4x + 3y = 3 b4x – 3y = 5 c3x – 4y = 3 d3x + 4y = 5